OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x->0 of -6+x/x^4

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

I'm not completely sure what this is asking for you to do.

OpenStudy (anonymous):

To find the limit,

OpenStudy (anonymous):

But i dont think its with direct substitution

OpenStudy (anonymous):

@Jim766

OpenStudy (anonymous):

@wio

OpenStudy (agent0smith):

Post it correctly, using parentheses, so we can tell what it is. Or post a screenshot

OpenStudy (anonymous):

Attempts?

OpenStudy (anonymous):

well I tried plugging in 0 for x, but oviously thats not going to work, :C

OpenStudy (anonymous):

So then i tried to expand the denominator

OpenStudy (anonymous):

Which still leaves 3 x's on the bottom

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

can you use L'Hopital's Rule ?

OpenStudy (anonymous):

I am not familiar with that

hartnn (hartnn):

ok, then factor out 'x' from the numerator, can you ?

OpenStudy (anonymous):

I've hear of the Power rule, but i'm in precalculus and apparently they dont show you that

OpenStudy (anonymous):

yes we have -3/x^3

OpenStudy (anonymous):

i mean -6/x^3

hartnn (hartnn):

how?? i was talking about this : \(\Large -6+x= x [\dfrac{-6}{x}+1]\) got this ?

OpenStudy (anonymous):

hm? How did you get that?

hartnn (hartnn):

nvm, then forget it..

hartnn (hartnn):

simply plug in x=0 in numerator and denominator, what do u get ?

OpenStudy (anonymous):

negative 6,

hartnn (hartnn):

and in denominator ?

OpenStudy (anonymous):

0 right,

OpenStudy (anonymous):

its -6+0/0^4

hartnn (hartnn):

yes, correct \(\Large \dfrac{-6}{0} = -\infty \) do you get why it is NEGATIVE infinity ?

OpenStudy (anonymous):

because the 6 is negative and the (asymptote) is going farther down negative y values?

OpenStudy (anonymous):

wait does that mean the limit doesnt exist?

hartnn (hartnn):

\(-\infty \) is the limiting value of your function. it is still a value. Limit will not exist ONLY IF , right hand limit does NOT equal to left hand limit.

OpenStudy (zarkon):

I would say the limit does not exist. It diverges to \(-\infty\)

OpenStudy (anonymous):

Ohhh, okay, thank you @hartnn and @Zarkon

hartnn (hartnn):

but left hand limit = right hand limit {= \(-\infty \)}so limit should exist. and welcome ^_^

OpenStudy (anonymous):

Ahh okay!! C:

OpenStudy (zarkon):

a limit converges if there is a number L such that the left and right hand limits give the same value L (L is some real number) infinity is not a real number

OpenStudy (anonymous):

I mean my answer choices are 6, 0, -6, and does not exist

OpenStudy (anonymous):

@Zarkon

hartnn (hartnn):

then you should surely choose DNE.

OpenStudy (anonymous):

oh okay, thank you C: